OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20 x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No Solution

OpenStudy (anonymous):

You need to solve two equations: (2x + 6) - 4 = 20 2x + 2 = 20 2x = 18 x = 9 AND -(2x + 6) - 4 = 20 -2x -6 - 4 = 20 -2x -10 = 20 -2x = 30 x = -15 Choice 3 is correct.

OpenStudy (anonymous):

thanks i found it difficult